OpenStudy (anonymous):

If f(x) = 8x2 − x3, find f'(2) and use it to find an equation of the tangent line to the curve y = 8x2 − x3 at the point (2, 24).

OpenStudy (anonymous):

f'(x)=16x-3x^2 plug in 2 and you get 20.

OpenStudy (anonymous):

now equation of a tangent line is y-y1=m(x-x1) so you have: y-24=20(x-2) that's it

OpenStudy (anonymous):

thanks!!!(: